Cos4A – Cos4B = 8(cosA-CosB)(CosA+CosB)(cosA-SinB)(CosA+SinB)

Grade:11

1 Answers

siddharth gupta
9 years ago
cos4A-cos4B=2cos22A-1-2cos22B+1=2(cos22A-cos22B)
                                                     =2((2cos2A-1)2-(2cos2B-1)2)
                                                      =2(4cos4A+1-4cos2A-4cos4B-1+4cos2B)
                                                      =8((cos4A-cos4B)-(cos2B-cos2A))
                                                      =8(cos2A-cos2B)((cos2A+cos2B)-1)
                                                      =8(cosA-cosB)(cosA+cosB)(cos2A-sin2B)
                                                                        =8(COSA-COSB)(COSA+COSB)(COSA-SINB)(COSA+SINB)
WITH REGARDS
SIDDHARTH PAVAN GUPTA